A comparative balance sheet for Martinez Corporation is presented as follows. Assets Cash Accounts receivable Inventory LandPrepare a statement of cash flows for 2020 for Martinez Corporation. (Show amounts that decrease cash flow with either a - siIssued common stock to retire $ of bonds outstandingDetermine Martinez Corporations current cash debt coverage, cash debt coverage, and free cash flow. (Round current cash debt
Answer #1
Cash Flow Statement
Indirect Method
Cash flow from Operating Activities
Net Income $          129,380
Adjustments
Depreciation $            26,900
Change In current assets & Liabilities
Increase in Accounts Receivable $           (15,900)
Decrease in Inventories $               9,100
Decrease in Accounts payable $           (13,100)
Total Adjustments $               7,000
Net Cash from operating activities $          136,380
Cash flow from Investing Activities
Sale of Land $            39,100
Purchase of Equipment $           (59,900)
Net Cash used in investing activities $          (20,800)
Cash flow from Financing Activities
Payment of cash dividends $           (64,680)
Net Cash used in financing activities $          (64,680)
Net Increase in cash $            50,900
Cash, beginning of the year $            22,000
Cash, ending of the year $            72,900
Non Cash Financing Activities
Issue of common stock to retire $50,000 of bonds outstanding

Current Cash Debt Coverage = Net Cash provided by Operating Activities / Average Current Liabilities
= $136380 / 42740 = 3.19

Cash Debt Coverage = Net Cash provided by Operating Activities / Average Total Liabilities
= $136380 / 217740 = 0.63

Free Cash FLow = Net Cash provided by Operating Activities - Capital Expenditure
= $136380 - 20800 = $115580

Martinez has high liquidity and its financial flexibility is strong

If you have any query, kindly comment with your query and please mark thumbs up.